### v2

@@ Abstract
Large Language Models (LLMs) have facilitated the definition of autonomous intelligent agents. Such agents have already demonstrated their potential in solving complex tasks in different domains. And they can further increase their performance when collaborating with other agents in a multi-agent system. However, the orchestration and coordination of these agents is still challenging, especially when they need to interact with humans as part of human-agentic collaborative workflows. These kinds of workflows need to be precisely specified so that it is clear whose responsible for each task, what strategies agents can follow to complete individual tasks or how decisions will be taken when different alternatives are proposed, among others. Current business process modeling languages fall short when it comes to specifying these new mixed collaborative scenarios. In this exploratory paper, we extend a well-known process modeling language (i.e., BPMN) to enable the definition of this new type of workflow. Our extension covers both the formalization of the new metamodeling concepts required and the proposal of a BPMN-like graphical notation to facilitate the definition of these workflows. Our extension has been implemented and is available as an open-source human-agentic workflow modeling editor on GitHub.
Keywords:
LLM Agent Collaborative Workflow BPMN Extension.
1 Introduction
@@
In our current information-rich society, the integration of agents, especially agents powered by Large Language Models (LLMs), is becoming more and more important to quickly perform many tasks. Agents can interact with the environment, make their own decisions, and learn from the received feedback. Moreover, often, agents do not work in isolation but as part of Multi-Agent Systems (MAS) [5] where agents cooperate (or compete) to achieve a common goal. This collaborative process is known as an agentic system. These systems have already demonstrated their superior performance against single-agent solutions [6].
@@
While agentic systems are performant in many tasks, complex scenarios require the participation of humans. Therefore, there is a need to precisely define this collaboration and how each participant interacts with each other. Unfortunately, on the one hand, current process modeling languages, such as BPMN, lack the modeling constructs to specify this collaboration between humans and agentic systems as new primitives to define the confidence of the agents, the strategies they can use to perform a task, or the process to reach a decision. And on the other hand, frameworks targeting the implementation of agentic workflows, such as LangGraph [^1] https://www.langchain.com/langgraph, minimize the participation of humans in the process and therefore are not expressive enough to model the human-agent interaction beyond very simple cases.
@@
Therefore, the goal of this paper is to enable the precise definition of human-agentic workflows. To this aim, we study how the Business Process Model and Notation (BPMN), one of the most well-known modeling languages for workflows, could be used to represent this new type of workflows, and then, based on the identified limitations, we propose a BPMN extension to enable their definition in BPMN. This extension has been implemented in an open source modeling tool available on GitHub.
@@
The rest of the paper is structured as follows. Section 2 provides the background and a running example. Section 3 shows the mapping of agentic concepts to BPMN, while Section 4 presents the extension to the BPMN to overcome the limitations found. Section 5 and  6 provide the extension definition and the proof of concept, respectively. Section 7 presents the related work. Finally, Section 8 concludes the paper and presents the roadmap.
2 Background & Running Example
@@
In this section, we briefly describe BPMN, the process modeling language we aim to extend; and present the main concepts of agentic systems. We end the section with a running example.
2.1 BPMN
@@
BPMN [10] has become the de-facto standard for business processes diagrams. It is defined by the Object Management Group (OMG) and specified as ISO standard (ISO/IEC 19510:2013). BPMN provides a metamodel and a notation to define and visualize business process models.
@@
2.2 Agentic systems
Although LLMs excel at multiple tasks [15], creating specialized instances of them to target specific tasks has shown promising results. These specialized LLMs are typically known as agents [11], leveraging the definition of intelligent agents [16]. LLM-based agents use the potential of LLMs, facilitating sophisticated interactions, decision-making, tool-use capabilities, and in-context learning through memory [6].
@@ 2.2 Agentic systems
2.3 Running Example
To illustrate our proposal, we will use a running example based on a simple resolution process for bug reports in a software project. The example process comprises five participants, two human and three agents.
The humans are a user, who reports the bug; and a maintainer, who reviews the final change proposal and resolves the bug.
The agents are responsible for solving the bug by implementing change proposals and deciding together the best option. There are three agents, one is used as a reviewer and the other two are specialized coding agents. Each agent comes with a level of uncertainty regarding the quality of all its actions. This value could be derived from the underlying LLM and/or the agent setup.
@@ 2.3 Running Example
3 Using BPMN to Model Human-Agentic Workflows
In this section, we study the current support of BPMN to model human-agentic workflows. The characteristics of agentic systems we aim at covering are the ones that concern the reflection and cooperation, while also addressing their non-deterministic behavior. Table 1 shows a possible mapping of BPMN elements to model human-agentic workflows. Note that this mapping is partial, as the current support of BPMN for this type of workflow is very limited, as we show in this section.
Table 1: Partial mapping of human-agentic workflow concepts to BPMN elements.
@@ 3 Using BPMN to Model Human-Agentic Workflows
To model the collaboration between agents, we consider two scenarios, depending on whether agents are located in lanes of the same or different pool/s. When located in the same pool, gateways could be used to route the flow towards multiple lanes, and groups could be used to set the collaboration strategy. If located in different pools, message flows could be used, but text annotations should be used to describe the strategy. For merging collaborations, in the former case we could use a complex gateway and specify the desired merging strategy as an annotation, while in the latter we could specify the strategy as an annotation in the incoming message flow. Note the need to use text annotations to describe the behavior in natural language.
@@
Representing human-agentic workflows in BPMN is challenging, as the standard BPMN does not provide specific elements to model agents or their interactions. In the running example, we cannot set uncertainty to determine the agents’ reliability or identify lanes or pools as agents. We can partially represent the collaboration and reflection strategies. We rely on natural language to describe the collaboration and reflection strategies, which can lead to ambiguity and misinterpretation. Furthermore, the intricate process of representing reflection can hinder the readability and understandability of the model. In the next section, we propose an extension to BPMN that introduces formalized syntax for defining collaboration and reflection strategies, rules for merging collaborative outcomes, and agents’ profiling (i.e., role and trustworthiness).
4 Extending BPMN to Model Human-Agentic Workflows
@@ 4.1 Agent Profiling
Agent profiling requires modeling (1) the role of the agents in collaboration scenarios, as it is how an LLM is initialized as an agent; and (2) the reliability of the agent’s output, or trust score, to address the non-deterministic nature of agents behavior. As shown in Section 3, in standard BPMN, representing agents as lanes or pools overlooks these aspects: (1) when having multiple lanes, each representing specialized agents, only lane names can be specified; and (2) there is no way to specify uncertainty for agent’s output.
@@ 4.3 Agent Collaboration
We model the CollaborationMode as the root of the hierarchy of collaboration modes, which can be either cooperation or competition, the former being the root of three types of cooperation (i.e., voting, role and debate). We extend both the inclusive and parallel gateways (see AgenticOR and AgenticAND, respectively), which also set the collaboration strategy (see CollaborationMode association). The inclusive gateway allows specifying a condition so that only specific agents are activated, while the parallel gateway diverges the flow to all its outputs, which is desired when all the participants from the diverged flow are required to collaborate Since agentic gateways are intended to represent collaboration, we only extend these two gateways. We do not consider the exclusive gateway, since the flow is allowed only to one output.
@@
The gateways allow the collaboration within pools, but if the set of agents is represented as another pool, message flows should be used (see MessageFlow). Like done for gateways, the outgoing message flow sets the collaboration strategy, while the incoming message flow sets the desired merging strategy.
@@
4.4 BPMN Notation Extension
According to the BPMN specification, the notation of an extension must not alter the notation of its elements, and must be as close as possible, in terms of look and feel, to it [10]. Our extension introduces four notations (see Table 2), following the graphical representation of the BPMN element being extended. To identify the extended elements, we use a marker representing an agent. The position of the marker is intended to not disturb the readability, and if possible, following the BPMN standards.
@@ 4.4 BPMN Notation Extension
The agentic task is represented with the agent marker in the top-left corner, as it is done with specific tasks (e.g., ManualTask). The reflection and collaboration strategies are indicated with a marker at the bottom of the shape. The reflection strategy is set as a letter inside the marker, where “s” stands for self-reflection, “c” for cross-reflection, and “h” for human-reflection. Note that the notation shown in Table 2 the “x” is used as a placeholder.
The agentic gateways are represented with the agent marker at the top-left side, without disturbing the shape of the gateway. The diverging gateway contains the collaboration marker in the bottom-right corner. The collaboration strategy is set as a letter below the marker, where “c” stands for competition, “d” for debate cooperation, “r” for role cooperation, and “v” for voting cooperation. The merging gateway must contain the merging marker in the bottom-right corner. The merging strategy is represented with two sets of letters below the marker, the first as the strategy class and the second as the merging strategy type. Thus, for voting strategies are “v-ma” for majority, “v-a” for absolute majority, and “v-mi” for minority; for role-based strategies are “r-l” for leader-driven and “r-c” for composed; while for competition strategies are “c-f” for fastest, and “c-mc” for most complete. As before, the notation for the diverging and merging gateway shows “x” as a placeholder.
@@
Table 2: Graphical notation of the extended elements.
Extension element Notation Extension element Notation
Agentic lane [Uncaptioned image] Agentic task [Uncaptioned image]
Diverging agentic gateway [Uncaptioned image] Merging agentic gateway [Uncaptioned image]
Outgoing agentic message flow [Uncaptioned image] Incoming agentic message flow [Uncaptioned image]
4.5 Using our Extension to Model the Running Example
@@
To represent the collaboration between agents, we use the agentic gateway, since the agents are represented as lanes of the same pool. The collaboration strategy applied is role cooperation, as denoted by the notation. After stating the collaboration strategy, the flow is divided towards the agents that collaborate. All the flows from this collaboration are merged into an agentic gateway following the strategy of leader-driven, as denoted by the notation. The remaining elements are compliant to standard BPMN.
5 Our Extension as a Standard BPMN Extension Definition
BPMN provides an extension mechanism to allow modeling domain-specific elements not included in the specification. In particular, BPMN 2.0 extension mechanism enables the approach of extension by addition, which consists of attaching new domain-specific elements to the predefined elements of the language.
Nevertheless, there is a lack of methodological guides to develop and publish specific extensions. Stroppi et al. [14] define a method (called BPMN+X) to transform a domain model into a BPMN-compliant extension by using UML profiles, which is the typical approach used in the UML world to define lightweight extensions [^2] The term lightweight extension is used to denote language extensions that are compatible with language semantics and that can be expressed using the own language extension mechanisms, enabling the direct use of the extension in any tool that supports the language. This is in contrast to heavyweight extensions that offer more complex extensions that enable richer semantics for the extension but require dedicated tooling support to the language. Given a BPMN extension defined as a profile, the BPMN+X method use mapping rules and automated model transformations to generate an XML Schema Extension Definition Document conforming to the official BPMN extension mechanism.
@@ 6 Proof of Concept
The extension has been implemented using Sirius  [^3] https://eclipse.dev/sirius/overview.html, an Eclipse project which allows you to easily create your own graphical modeling workbench by leveraging Eclipse Modeling technologies such as EMF and GMF. Aconite [12], an open-source tool based on a textual notation to reduce manual work to produce Sirius-based graphical notations, has been used to automatically generate the Sirius-based implementation. Figure 8 shows a screenshot of the tool.
The tool repository [^4] https://github.com/BESSER-PEARL/agentic-bpmn includes the files that use the Aconite annotations and the examples illustrated in this paper. The tool includes a main view of the diagram and a palette to allow the user to drag and drop the elements into the diagram. The palette contains the basic representation of BPMN, plus the extension elements proposed in this paper.
@@ 7 Related Work
We classify the related work in two groups: proposals that try to model agents in general process modeling languages and agent-specific tools that may include a language to orchestrate agentic workflows.
@@
As an example of the first group we have Küster et al. [8] that combine agent-oriented software engineering with business process design using BPMN. Nevertheless, they do not extend BPMN and therefore suffer from the limitations we stated in Section 3. Endert et al. [3] propose a mapping of agents to business processes, in particular agents that follow the belief-desire-intention (BDI) paradigm. However, while this proposal provides ways of mapping agent-specific concepts following the BDI paradigm to BPMN, they do not cover advanced aspects of the current generation of agents such as the reflection strategies.
@@
Other BPMN approaches partially overlap with our uncertainty definition. Ceballos et al. [1] propose a BPMN Business Process Diagram (BPD) normal form based on Activity Theory [4] that can be used for representing the dynamics of a collective human activity from the perspective of a subject. Herbert and Sharp [7] proposed a BPMN extension introducing the uncertainty in sequence flows. The extension includes probabilistic flows and rewards associated to the execution of tasks. Note that the uncertainty is determined in the sequence flows rather than in the participant or activity BPMN element. Our proposal, in the other hand, introduces the uncertainty at the participant level, then propagating to all the other elements. Both types of uncertainty could be combined.
@@
Regarding specific tools to create agentic systems, we have a couple that include a graphical modeling of the process. Two representative examples are LangGraph [^5] https://www.langchain.com/langgraph that allows modeling the action flow of an agent (or a set of agents) using cyclic graphs and Flowise [^6] https://github.com/FlowiseAI/Flowise, an open-source low-code tool for orchestrating LLM agents. In both cases, the workflows are purely focused on agent collaborations. Humans can only trigger the actions but are not supposed to collaborate with the agent to accomplish them. In contrast, our approach enables a human-agentic collaboration.
The scenarios presented by these proposals can also be covered by our extension, leveraging on standard BPMN concepts such as sequence flow elements for hierarchical or sequential flows, or lanes and pools to represent single and multiple agents, along with the extension elements we proposed (e.g., collaboration strategies). Our extension also enables representing humans not only as validators, but to define complex actions with or without interaction with agents.
8 Conclusions and Further Work

### v3

@@ Abstract
Large Language Models (LLMs) have facilitated the definition of autonomous intelligent agents. Such agents have already demonstrated their potential in solving complex tasks in different domains. And they can further increase their performance when collaborating with other agents in a multi-agent system. However, the orchestration and coordination of these agents is still challenging, especially when they need to interact with humans as part of human-agentic collaborative workflows. These kinds of workflows need to be precisely specified so that it is clear whose responsible for each task, what strategies agents can follow to complete individual tasks or how decisions will be taken when different alternatives are proposed, among others. Current business process modeling languages fall short when it comes to specifying these new mixed collaborative scenarios. In this exploratory paper, we extend a well-known process modeling language (i.e., BPMN) to enable the definition of this new type of workflow. Our extension covers both the formalization of the new modeling concepts required and the proposal of a BPMN-like graphical notation to facilitate the definition of these workflows. Our extension has been implemented and is available as an open-source human-agentic workflow modeling editor on GitHub.
I Introduction
@@
In our current information-rich society, the integration of agents, especially agents powered by Large Language Models (LLMs), is becoming more and more important to quickly perform many tasks [17]. Agents can interact with the environment, make their own decisions, and learn from the received feedback. Moreover, often, agents do not work in isolation but as part of Multi-Agent Systems (MAS) where agents cooperate (or compete) to achieve a common goal [6]. This collaborative process is known as an agentic system. These systems have already demonstrated their superior performance against single-agent solutions [6].
@@
While agentic systems are performant in many tasks, complex scenarios require the participation of humans [16]. Therefore, there is a need to precisely define this collaboration and how each participant interacts with each other. Unfortunately, we argue that current process modeling languages, such as BPMN, lack the modeling constructs to specify this collaboration between humans and agentic systems as new primitives to define the confidence of the agents, the strategies they can use to perform a task, or the process to reach a decision. Furthermore, frameworks targeting the implementation of agentic workflows, such as LangGraph [^1] https://www.langchain.com/langgraph, minimize the participation of humans in the process and therefore are not expressive enough to model the human-agent interaction beyond very simple cases.
@@
The goal of this paper is to enable the precise definition of human-agentic workflows. To this aim, we study how the Business Process Model and Notation (BPMN), one of the most well-known modeling languages for workflows, could be used to represent this new type of workflow, and then, based on the identified limitations, we propose a BPMN extension to enable their definition in BPMN. Note that our approach can serve as a blueprint for extending other workflow languages, as the conceptual elements we identify are largely notation-independent. This extension has been implemented in an open source modeling tool available on GitHub.
@@
The rest of the paper is structured as follows. Section II provides the background and a running example. Section III shows the mapping of agentic concepts to BPMN, while Section IV presents the extension to the BPMN to overcome the limitations found. Section V and  VI provide the extension definition and the proof of concept, respectively. Section VII presents the related work. Finally, Section VIII concludes the paper and presents the roadmap.
II Background & Running Example
@@
In this section, we briefly describe BPMN, the language we aim to extend; and present the main concepts of agentic systems. We end the section with a running example.
II-A BPMN
@@
BPMN [11] has become the de-facto standard for business processes diagrams. It is defined by the Object Management Group (OMG) and specified as ISO standard (ISO/IEC 19510:2013). BPMN provides a metamodel and a notation to define and visualize business process models. It provides an extension mechanism to allow modeling domain-specific elements not included in the specification. In particular, BPMN 2.0 extension mechanism enables the approach of extension by addition, which consists of attaching new domain-specific elements to the predefined elements of the language.
@@
II-B Agentic systems
LLMs excel at multiple tasks [15], but creating specialized LLMs instances to target specific tasks has shown promising results. These specialized LLMs are typically known as agents [17]. LLM-based agents use the potential of LLMs, facilitating sophisticated interactions, decision-making, tool-use capabilities, and in-context learning through memory [6].
@@ II-B Agentic systems
II-C Running Example
To illustrate our proposal, we will use a running example based on a simple resolution process for bug reports in a software project. The example process comprises five participants, two humans and three agents.
The humans are a user, who reports the bug; and a maintainer, who reviews the final change proposal and resolves the bug. The agents are responsible for solving the bug by implementing change proposals and deciding together the best option. There are three agents, one is used as a reviewer and the other two are specialized coding agents. Each agent comes with a level of uncertainty regarding the quality of all its actions. This value could be derived from the underlying LLM and/or the agent setup.
@@ II-C Running Example
III Using BPMN to Model Human-Agentic Workflows
In this section, we study the current support of BPMN to model human-agentic workflows. The characteristics of agentic systems we aim at covering are the ones that concern the reflection and cooperation, while also addressing their non-deterministic behavior. Table I shows a possible mapping of BPMN elements to model human-agentic workflows. Note that this mapping is partial, as the current support of BPMN for this type of workflow is very limited, as we show in this section.
TABLE I: Partial mapping of human-agentic workflow concepts to BPMN elements.
@@ III Using BPMN to Model Human-Agentic Workflows
To model the collaboration between agents, we consider two scenarios, depending on whether agents are located in lanes of the same or different pool/s (i.e., collaboration diagram). When located in the same pool, gateways could be used to route the flow towards multiple lanes, and groups could be utilized to set the collaboration strategy. If located in different pools, message flows could be used, but text annotations should be added to describe the strategy. For merging collaborations, in the first scenario we could use a complex gateway and specify the desired merging strategy as an annotation, while in the second scenario we could specify the strategy as an annotation in the incoming message flow. Note the need to use text annotations to describe the behavior. BPMN also offers additional collaboration mechanisms such as choreographies (formalized coordination of interactions between participants), but these still require supplementary descriptions for agent-specific behaviors since their focus is on the exchange of information rather than the orchestration of their work.
@@
Representing human-agentic workflows in BPMN is challenging, as the standard BPMN does not provide specific elements to model agents or their interactions. In the running example, we cannot set uncertainty to determine the agents’ reliability or identify lanes or pools as agents. We can partially represent the collaboration and reflection strategies. We rely on natural language to describe the collaboration and reflection strategies, which can lead to ambiguity and misinterpretation. Furthermore, the intricate process of representing reflection can hinder the readability and understandability of the model.
BPMN was chosen for several key reasons: (1) widespread adoption, BPMN is the de facto standard for business process modeling with broad industrial and academic support; (2) expressivity, BPMN already provides comprehensive constructs for modeling workflows involving human participants; (3) extensibility, BPMN offers an extension mechanism that enables domain-specific additions while maintaining compliance with the standard; (4) execution capabilities, BPMN models can be directly executed by business process engines, making the transition from conceptual models to implementation easier; and (5) integration with existing processes, many organizations already use BPMN, making it easier to incorporate agent-based elements into existing business processes.
Therefore, despite the identified limitations, we believe that BPMN still provides a good starting point compared to alternative modeling languages. While languages like UML Activity Diagrams or YAWL could potentially be extended in similar ways, BPMN’s widespread usage in industries means that practitioners are already familiar with its core concepts, reducing the learning curve for our extended notation.
In the next section, we propose an extension to BPMN that introduces formalized syntax for defining collaboration and reflection strategies, rules for merging collaborative outcomes, and agents’ profiling (i.e., role and trustworthiness).
IV Extending BPMN to Model Human-Agentic Workflows
@@ IV-A Agent Profiling
Agent profiling requires modeling (1) the role of the agents in collaboration scenarios, as it is how an LLM is initialized as an agent; and (2) the reliability of the agent’s output, or trust score, to address the non-deterministic nature of agents behavior. As shown in Section III, in standard BPMN, representing agents as lanes or pools overlooks these aspects: (1) when having multiple lanes, each representing specialized agents, only lane names can be specified; and (2) there is no way to specify uncertainty for agent’s output.
@@ IV-C Agent Collaboration
We model the CollaborationMode as the root of the hierarchy of collaboration modes, which can be either cooperation or competition, the former being the root of three types of cooperation (i.e., voting, role and debate). We extend both the inclusive and parallel gateways (see AgenticOR and AgenticAND, respectively), which also set the collaboration strategy (see CollaborationMode association). The inclusive gateway allows specifying a condition so that only specific agents are activated, while the parallel gateway diverges the flow to all its outputs, which is desired when all the participants from the diverged flow are required to collaborate. Since agentic gateways are intended to represent collaboration, we only extend these two gateways. We do not consider the exclusive gateway, since the flow is allowed only to one output.
@@
The gateways allow the collaboration within pools, but if the set of agents is represented as another pool, message flows should be used (see MessageFlow). We extend the message flow, rather than choreography tasks, since it allows for an elementary depiction of the collaboration. Like done for gateways, the outgoing message flow sets the collaboration strategy, while the incoming message flow sets the desired merging strategy.
@@
IV-D BPMN Notation Extension
According to the BPMN specification, extensions notation must not alter the notation of its elements, and must be as close as possible, in terms of look and feel, to it [11]. Our extension introduces four new elements (see Table II) with a graphical representation close the BPMN element being extended.
When choosing these elements, we also aimed to stick to the principles for effective visual notations by Moody [10]. For instance, to be compliant with the semiotic clarity principle, we maintain a 1:1 correspondence between semantic constructs and graphical symbols, with each agentic concept (lane, task, gateway, message flow) having its own distinct visual representation (see Table II).
To identify the extended elements, we use a marker representing an agent. Our agent marker follows the perceptual discriminability principle, since it clearly distinguishes agentic elements from standard BPMN elements, while maintaining the basic shape of the original BPMN symbols to preserve familiarity. Furthermore, the agent icon intuitively suggests intelligence and automation being semantically immediate, following the semantic transparency principle.
On the other hand, we minimize the introduction of new symbols by using a consistent agent marker with letter modifiers rather than creating entirely new shapes, following the graph economy and dual coding principle (i.e., use both graphical elements and textual annotations to enhance cognition).
@@ IV-D BPMN Notation Extension
The agentic task is represented with the agent marker in the top-left corner, as it is done with specific tasks (e.g., ManualTask). The reflection and collaboration strategies are indicated with a marker at the bottom of the shape. The reflection strategy is set as a letter inside the marker, where “s” stands for self-reflection, “c” for cross-reflection, and “h” for human-reflection. Note that the notation shown in Table II the “x” is used as a placeholder.
The agentic gateways are represented with the agent marker at the top-left side, without disturbing the shape of the gateway. The diverging gateway contains the collaboration marker in the bottom-right corner. This way the visual distance between symbols is greater [10]. The same principle has been applied for the remaining notations. The collaboration strategy is set as a letter below the marker, where “c” stands for competition, “d” for debate cooperation, “r” for role cooperation, and “v” for voting cooperation. The merging gateway must contain the merging marker in the bottom-right corner. The merging strategy is represented with two sets of letters below the marker, the first as the strategy class and the second as the merging strategy type. Thus, for voting strategies are “v-ma” for majority, “v-a” for absolute majority, and “v-mi” for minority; for role-based strategies are “r-l” for leader-driven and “r-c” for composed; while for competition strategies are “c-f” for fastest, and “c-mc” for most complete.
@@
TABLE II: Graphical notation of the extended elements.
Extension element Notation
Agentic lane [Uncaptioned image]
Agentic task [Uncaptioned image]
Diverging agentic gateway [Uncaptioned image]
Merging agentic gateway [Uncaptioned image]
Outgoing agentic message flow [Uncaptioned image]
Incoming agentic message flow [Uncaptioned image]
IV-E Using our Extension to Model the Running Example
@@
To represent the collaboration between agents, we use the agentic gateway, since agents are represented as lanes of the same pool. The collaboration strategy applied is role cooperation, as denoted by the notation. After stating the collaboration strategy, the flow is divided towards the agents that collaborate. All flows from this collaboration are merged into an agentic gateway following the leader-driven, strategy as denoted by the notation. The remaining elements are compliant to standard BPMN.
V Our Extension as a Standard BPMN Extension Definition
BPMN provides an extension mechanism to allow modeling domain-specific elements not included in the specification.
Nevertheless, there is a lack of methodological guides to develop and publish specific extensions. Stroppi et al. [14] define a method (called BPMN+X) to transform a domain model into a BPMN-compliant extension by using UML profiles, which is the typical approach used in the UML world to define lightweight extensions [^2] The term lightweight extension is used to denote language extensions that are compatible with language semantics and that can be expressed using the own language extension mechanisms, enabling the direct use of the extension in any tool that supports the language. This is in contrast to heavyweight extensions that offer more complex extensions that enable richer semantics for the extension but require dedicated tooling support to the language. Given a BPMN extension defined as a profile, the BPMN+X method uses mapping rules and automated model transformations to generate an XML Schema Extension Definition Document conforming to the official BPMN extension mechanism.
@@ VI Proof of Concept
The extension has been implemented using Sirius  [^3] https://eclipse.dev/sirius/overview.html, an Eclipse project which allows you to easily create your own graphical modeling workbench by leveraging Eclipse Modeling technologies such as EMF and GMF. Aconite [12], a tool that helps produce Sirius-based graphical notations, has been used to automatically generate the Sirius-based implementation. Figure 8 shows a screenshot.
The tool repository [^4] https://github.com/BESSER-PEARL/agentic-bpmn includes the files that use the Aconite annotations and the examples illustrated in this paper. The tool includes a main view of the diagram and a palette to allow the user to drag and drop the elements into the diagram. The palette contains the basic representation of BPMN, plus the extension elements proposed in this paper. Furthermore, the repository also includes several examples from the literature to illustrate the use of the extension.
@@ VII Related Work
We classify the related work into two groups: proposals that try to model agents in general process modeling languages and agent-specific tools that may include a language to orchestrate agentic workflows.
@@
In the first group we have Küster et al. [8] that combine agent-oriented software engineering with business process design. Nevertheless, they do not extend BPMN and therefore suffer from the limitations stated in Section III. Endert et al. [4] propose a mapping of BDI agents to business processes. However, while this proposal provides ways of mapping agent-specific concepts following the BDI paradigm to BPMN, they do not cover advanced aspects of the current generation of agents such as reflection strategies.
@@
Other BPMN approaches partially cover uncertainty concerns. Ceballos et al. [2] propose a BPMN Business Process Diagram (BPD) normal form based on Activity Theory [5] that can be used for representing the dynamics of a collective human activity from the perspective of a subject. Herbert and Sharp [7] proposed a BPMN extension introducing the uncertainty in sequence flows. The extension includes probabilistic flows and rewards associated to the execution of tasks. Note that the uncertainty is determined in the sequence flows rather than in the participant or activity BPMN element. Our proposal introduces uncertainty at the participant level, then propagating to all the other elements. Both types of uncertainty could be combined.
@@
Regarding specific tools to create agentic systems, a couple include the graphical modeling of the process. Two representative examples are LangGraph [^5] https://www.langchain.com/langgraph that allows modeling the action flow of an agent (or a set of agents) using cyclic graphs and Flowise [^6] https://github.com/FlowiseAI/Flowise, an open-source low-code tool for orchestrating LLM agents. In both cases, the workflows are purely focused on agent collaborations. Humans can only trigger the actions but are not supposed to collaborate with the agent to accomplish them. In contrast, our extension considers humans as active participants and enables representing complex human-agent interactions.
VIII Conclusions and Further Work